# for Large Language Model Compression

### v1

@@ 3.2 Algorithms
Let $\hat{{{\bm{W}}}}_{U}$ be searched over the matrix multiplication form ${{\bm{W}}}_{U}{\bm{S}}_{k}$, where ${\bm{S}}_{k}$ is a $k$-column selection matrix, and let $\hat{{\bm{W}}}_{D}$ be searched over $\mathbb{R}^{k\times d_{h}}$. The optimal $\hat{{\bm{W}}}^{*}_{D}$ is then given by: $({\bm{S}}_{k}^{\top}{\bm{C}}^{1/2}_{\sigma}{\bm{S}}_{k})^{\dagger}{\bm{S}}_{k}^{\top}{\bm{C}}^{1/2}_{\sigma}{\bm{W}}_{D}$. Using ${{\bm{W}}}_{U}{\bm{S}}_{k}$ and $\hat{{\bm{W}}}^{*}_{D}$ as the compressed matrices, the Type-I reconstruction error in equation 6 satisfies:
where $\mathcal{E}_{\text{Nys}}({\bm{C}}_{\sigma}^{\frac{1}{2}})$ denotes the Nyström approximation error, defined in Def. 1, relative to the activation correlation ${\bm{C}}_{\sigma}\triangleq\sum_{i=1}^{N}\sigma({\bm{X}}_{i}{{\bm{W}}_{U}})^{\top}\sigma({\bm{X}}_{i}{{\bm{W}}_{U}})$, using the same ${\bm{S}}_{k}$ in the compression of ${{{\bm{W}}}}_{U}$.
@@
Suppose that the rank $k$ and the scores $s_{i}$ in Algorithm 1 are chosen such that there exists an error $\varepsilon>0$ satisfying $\varepsilon\geq\sum_{i=k+1}^{d_{int}}s_{i}$, then the Type-I modular reconstruction error in equation 6 is bounded by $V_{\text{I}}\leq\|{\bm{W}}_{D}\|_{2}^{2}\frac{\varepsilon^{2}{d^{2}_{int}}}{k^{2}(1-\varepsilon)^{2}}\sum_{i=k+1}^{d_{int}}\sigma_{i}({\bm{C}}_{\sigma})$, where $d_{\text{int}}$ and $\sigma_{i}$ denote the intermediate dimension (i.e., the input dimension of $W_{D}$) and singular values, respectively.
@@
2:$s_{i}$ $\leftarrow[{\bm{C}}^{1/2}_{\sigma}({\bm{C}}^{1/2}_{\sigma}+\lambda{\bm{I}})^{-1}]_{ii}$, for $i=1,\dots,d_{\text{int}}$triangleright Calculate the ridge leverage score
@@
4:return $({{\bm{W}}}_{U},{{{\bm{W}}}}_{D})\leftarrow({{\bm{W}}}_{U}{\bm{S}}_{k},\leavevmode\nobreak\ \leavevmode\nobreak\ ({\bm{S}}_{k}^{\top}{\bm{C}}^{1/2}_{\sigma}{\bm{S}}_{k})^{\dagger}{\bm{S}}_{k}^{\top}{\bm{C}}^{1/2}_{\sigma}{\bm{W}}_{D})$
@@ 4.3 Zero-shot performance
0% Dense 5.68 8.35 73% 42% 79% 70% 50% 27% 34% 54%
20% FWSVD [27] 1727 2152 31% 23% 56% 50% 26% 21% 15% 32%
ASVD [71] 11.14 16.55 53% 27% 68% 64% 41% 24% 25% 43%
SVD-LLM [62] 7.94 16.22 58% 29% 69% 58% 43% 24% 22% 44%
MoDeGPT (ours) 6.53 39.17 70% 36% 74% 69% 50% 26% 31% 51%
40% FWSVD [27] 18156 20990 26% 22% 53% 51% 26% 21% 16% 30%
ASVD [71] 1407 3292 28% 22% 55% 48% 26% 19% 12% 30%
SVD-LLM [62] 13.11 63.75 42% 25% 60% 58% 33% 21% 19% 37%
MoDeGPT (ours) 9.39 60.55 58% 30% 65% 64% 40% 23% 22% 43%
@@ A.1 Proof of Theorem 1 and Proposition 1: MLP Compression with Nyström Approximation
Let $\hat{{{\bm{W}}}}_{U}$ be searched over the matrix multiplication form ${{\bm{W}}}_{U}{\bm{S}}_{k}$, where ${\bm{S}}_{k}$ is a $k$-column selection matrix, and let $\hat{{\bm{W}}}_{D}$ be searched over $\mathbb{R}^{k\times d_{h}}$. The optimal $\hat{{\bm{W}}}^{*}_{D}$ can then be expressed as: $({\bm{S}}_{k}^{\top}{\bm{C}}^{\frac{1}{2}}_{\sigma}{\bm{S}}_{k})^{\dagger}{\bm{S}}_{k}^{\top}{\bm{C}}^{\frac{1}{2}}_{\sigma}{\bm{W}}_{D}$. Using ${{\bm{W}}}_{U}{\bm{S}}_{k}$ and $\hat{{\bm{W}}}^{*}_{D}$ as the compressed matrices, the Type-I reconstruction error in equation 6 satisfies:
@@
where ${\bm{C}}_{\sigma}$ is the empirical correlation matrix of latent features ${\bm{C}}_{\sigma}=\sum_{i=1}^{N}\sigma({\bm{X}}_{i}{\bm{W}}_{U})^{\top}\sigma({\bm{X}}_{i}{\bm{W}}_{U})$, and (a) follows from Assumption 1. Setting the gradient of the last expression with respect to $\hat{{\bm{W}}}_{D}$ to zero, we obtain the optimal down matrix $\hat{{\bm{W}}}^{*}_{D}=\left({\bm{S}}_{k}^{\top}{\bm{C}}^{\frac{1}{2}}_{\sigma}{\bm{S}}_{k}\right)^{\dagger}{\bm{S}}_{k}^{\top}{\bm{C}}^{\frac{1}{2}}_{\sigma}{\bm{W}}_{D}$. After Plugging this back to the objective, we can further simplify the objective into,
$\min_{{\bm{S}}_{k}}\left\|\left({\bm{C}}^{\frac{1}{2}}_{\sigma}-{\bm{C}}^{\frac{1}{2}}_{\sigma}{\bm{S}}_{k}\left({\bm{S}}_{k}^{\top}{\bm{C}}^{\frac{1}{2}}_{\sigma}{\bm{S}}_{k}\right)^{\dagger}{\bm{S}}_{k}^{\top}{\bm{C}}^{\frac{1}{2}}_{\sigma}\right){\bm{W}}_{D}\right\|_{F}^{2}$
@@
Now, observe that the error on the right side of equation 14 is equal to the Nyström matrix approximation to the matrix ${\bm{C}}^{\frac{1}{2}}_{\sigma}$ in Definition 1. Hence, the variable ${\bm{S}}_{k}$ can be optimized with any Nyström approximation algorithm [21]. In this work, we adapt a deterministic Nyström algorithm, Algorithm 1, that has the theoretical guarantee proved in the next proposition. ∎
@@
Suppose that the rank $k$ and the scores $s_{i}$ in Algorithm 1 are chosen such that there exists an error $\varepsilon>0$ satisfying $\varepsilon\geq\sum_{i=k+1}^{d_{int}}s_{i}$, then the Type-I modular reconstruction error in equation 6 is bounded by $V_{I}\leq\|{\bm{W}}_{D}\|_{2}^{2}\frac{\varepsilon^{2}{d^{2}_{int}}}{k^{2}(1-\varepsilon)^{2}}\sum_{i=k+1}^{d_{int}}\sigma_{i}({\bm{C}}_{\sigma})$, where $d_{\text{int}}$ and $\sigma_{i}$ denote the intermediate dimension (i.e., the input dimension of $W_{D}$) and singular values, respectively.
@@
Since our column selection is equivalent to applying the deterministic ridge leverage score sampling (DRLS) to ${\bm{C}}_{\sigma}^{\frac{1}{4}}$ [38] , Theorem 1 in [38] implies that
$(1-\varepsilon){\bm{C}}^{\frac{1}{2}}_{\sigma}-\frac{\epsilon}{k}\left\|({\bm{C}}_{\sigma}^{\frac{1}{4}})_{\backslash k}\right\|_{F}^{2}\mathbf{I}\preceq{\bm{C}}_{\sigma}^{\frac{1}{4}}{\bm{S}}_{k}{\bm{S}}_{k}^{\top}{\bm{C}}_{\sigma}^{\frac{1}{4}}\preceq{\bm{C}}_{\sigma}^{\frac{1}{2}}$ (15)
@@
Next, we define ${\bm{P}}={\bm{C}}^{\frac{1}{4}}_{\sigma}{\bm{S}}_{k}({\bm{S}}_{k}^{\top}{\bm{C}}^{\frac{1}{2}}_{\sigma}{\bm{S}}_{k})^{\dagger}{\bm{S}}_{k}^{\top}{\bm{C}}^{\frac{1}{4}}_{\sigma}$. We note that ${\bm{P}}$ is the projection matrix of the column space of ${\bm{C}}^{\frac{1}{4}}_{\sigma}{\bm{S}}$. Now, we multiply ${\bm{I}}-{\bm{P}}$ to both sides in the previous inequality to get
$({\bm{I}}-{\bm{P}}){\bm{C}}^{\frac{1}{2}}_{\sigma}({\bm{I}}-{\bm{P}})$ (17)
@@
where in the last inequality we use the fact that ${\bm{I}}-{\bm{P}}\preceq{\bm{I}}$ and that ${\bm{I}}-{\bm{P}}$ is the orthogonal projection to the orthogonal complement of the column space ${\bm{C}}_{\sigma}^{\frac{1}{4}}{\bm{S}}$ so that ${\bm{S}}^{\top}{\bm{C}}_{\sigma}^{\frac{1}{4}}({\bm{I}}-{\bm{P}})=\mathbf{0}$. Now, we have
$\|({\bm{I}}-{\bm{P}}){\bm{C}}^{\frac{1}{4}}_{\sigma}{\bm{C}}^{\frac{1}{4}}_{\sigma}({\bm{I}}-{\bm{P}})\|_{2}\leq\frac{\varepsilon}{k(1-\varepsilon)}\left\|({\bm{C}}_{\sigma}^{\frac{1}{4}})_{\backslash k}\right\|_{F}^{2}=\frac{\varepsilon}{k(1-\varepsilon)}\sum_{i=k+1}^{d_{\text{int}}}\sigma_{i}({\bm{C}}^{\frac{1}{2}}_{\sigma})$ (20)
@@
Since ${\bm{C}}^{\frac{1}{4}}_{\sigma}{\bm{P}}{\bm{C}}^{\frac{1}{4}}_{\sigma}={\bm{C}}^{\frac{1}{2}}_{\sigma}{\bm{S}}_{k}\left({\bm{S}}_{k}^{\top}{\bm{C}}^{\frac{1}{2}}_{\sigma}{\bm{S}}_{k}\right)^{\dagger}{\bm{S}}_{k}^{\top}{\bm{C}}^{\frac{1}{2}}_{\sigma}$, the inequality is equivalent to
$\|{\bm{C}}^{\frac{1}{2}}_{\sigma}-{\bm{C}}^{\frac{1}{2}}_{\sigma}{\bm{S}}\left({\bm{S}}^{\top}{\bm{C}}^{\frac{1}{2}}_{\sigma}{\bm{S}}\right)^{\dagger}{\bm{S}}^{\top}{\bm{C}}^{\frac{1}{2}}_{\sigma}\|_{2}\leq\frac{\varepsilon}{k(1-\varepsilon)}\sum_{i=k+1}^{d_{\text{int}}}\sigma_{i}({\bm{C}}^{\frac{1}{2}}_{\sigma}).$ (22)
@@
$\mathcal{E}_{\text{Nys}}^{2}({\bm{C}}^{\frac{1}{2}}_{\sigma})\leavevmode\nobreak$ $\overset{(a)}{\leq}{d_{\text{int}}}\|{\bm{C}}^{\frac{1}{2}}_{\sigma}-{\bm{C}}^{\frac{1}{2}}_{\sigma}{\bm{S}}\left({\bm{S}}^{\top}{\bm{C}}^{\frac{1}{2}}_{\sigma}{\bm{S}}\right)^{\dagger}{\bm{S}}^{\top}{\bm{C}}^{\frac{1}{2}}_{\sigma}\|^{2}_{2}$
$\overset{(b)}{\leq}\frac{\varepsilon^{2}{d_{int}}}{k^{2}(1-\varepsilon)^{2}}\left(\sum_{i=k+1}^{d_{int}}\sigma_{i}({\bm{C}}^{\frac{1}{2}}_{\sigma})\right)^{2}$
$\overset{(c)}{\leq}\frac{\varepsilon^{2}{d^{2}_{int}}}{k^{2}(1-\varepsilon)^{2}}\sum_{i=k+1}^{d_{int}}\sigma_{i}({\bm{C}}_{\sigma}),$ (23)

### v3

@@ 3.2 Algorithms
Let $\hat{{{\bm{W}}}}_{U}$ be searched over the matrix multiplication form ${{\bm{W}}}_{U}{\bm{S}}_{k}$, where ${\bm{S}}_{k}$ is a $k$-column selection matrix, and let $\hat{{\bm{W}}}_{D}$ be searched over $\mathbb{R}^{k\times d_{h}}$. The optimal $\hat{{\bm{W}}}^{*}_{D}$ is then given by: $({\bm{S}}_{k}^{\top}{\bm{C}}_{\sigma}{\bm{S}}_{k})^{\dagger}{\bm{S}}_{k}^{\top}{\bm{C}}_{\sigma}{\bm{W}}_{D}$. Using ${{\bm{W}}}_{U}{\bm{S}}_{k}$ and $\hat{{\bm{W}}}^{*}_{D}$ as the compressed matrices, the Type-I reconstruction error in equation 6 satisfies:
$V_{\text{I}}\leq\left\|{\bm{W}}_{D}\right\|_{2}^{2}\|{\bm{C}}^{-1}_{\sigma}\|_{2}\mathcal{E}^{2}_{\text{Nys}}({\bm{C}}_{\sigma}),$ (7)
where $\mathcal{E}_{\text{Nys}}({\bm{C}}_{\sigma})$ denotes the Nyström approximation error, defined in Def. 1, relative to the activation correlation ${\bm{C}}_{\sigma}\triangleq\sum_{i=1}^{N}\sigma({\bm{X}}_{i}{{\bm{W}}_{U}})^{\top}\sigma({\bm{X}}_{i}{{\bm{W}}_{U}})$, using the same ${\bm{S}}_{k}$ in the compression of ${{{\bm{W}}}}_{U}$.
@@
Suppose that the rank $k$ and the scores $s_{i}$ in Algorithm 1 are chosen such that there exists an error $\varepsilon>0$ satisfying $\varepsilon\geq\sum_{i=k+1}^{d_{int}}s_{i}$, then the Type-I modular reconstruction error in equation 6 is bounded by $V_{I}\leq\|{\bm{W}}_{D}\|_{2}^{2}\|{\bm{C}}^{-1}_{\sigma}\|_{2}\frac{\varepsilon^{2}{d^{2}_{int}}}{k^{2}(1-\varepsilon)^{2}}\sum_{i=k+1}^{d_{int}}\sigma_{i}^{2}({\bm{C}}_{\sigma})$, where $d_{\text{int}}$ and $\sigma_{i}$ denote the intermediate dimension (i.e., the input dimension of $W_{D}$) and singular values, respectively.
@@
2:$s_{i}$ $\leftarrow[{\bm{C}}_{\sigma}({\bm{C}}_{\sigma}+\lambda{\bm{I}})^{-1}]_{ii}$, for $i=1,\dots,d_{\text{int}}$triangleright Calculate the ridge leverage score
@@
4:return $({{\bm{W}}}_{U},{{{\bm{W}}}}_{D})\leftarrow({{\bm{W}}}_{U}{\bm{S}}_{k},\leavevmode\nobreak\ \leavevmode\nobreak\ ({\bm{S}}_{k}^{\top}{\bm{C}}_{\sigma}{\bm{S}}_{k})^{\dagger}{\bm{S}}_{k}^{\top}{\bm{C}}_{\sigma}{\bm{W}}_{D})$
@@ 4.3 Zero-shot performance
0% Dense 5.68 8.35 73 42 79 70 50 27 34 54
20% FWSVD [27] 1727 2152 31 23 56 50 26 21 15 32
ASVD [71] 11.14 16.55 53 27 68 64 41 24 25 43
SVD-LLM [62] 7.94 16.22 58 29 69 58 43 24 22 44
MoDeGPT (ours) 6.53 39.17 70 36 74 69 50 26 31 51
40% FWSVD [27] 18156 20990 26 22 53 51 26 21 16 30
ASVD [71] 1407 3292 28 22 55 48 26 19 12 30
SVD-LLM [62] 13.11 63.75 42 25 60 58 33 21 19 37
MoDeGPT (ours) 9.39 60.55 58 30 65 64 40 23 22 43
@@ A.1 Proof of Theorem 1 and Proposition 1: MLP Compression with Nyström Approximation
Let $\hat{{{\bm{W}}}}_{U}$ be searched over the matrix multiplication form ${{\bm{W}}}_{U}{\bm{S}}_{k}$, where ${\bm{S}}_{k}$ is a $k$-column selection matrix, and let $\hat{{\bm{W}}}_{D}$ be searched over $\mathbb{R}^{k\times d_{h}}$. The optimal $\hat{{\bm{W}}}^{*}_{D}$ can then be expressed as: $({\bm{S}}_{k}^{\top}{\bm{C}}_{\sigma}{\bm{S}}_{k})^{\dagger}{\bm{S}}_{k}^{\top}{\bm{C}}_{\sigma}{\bm{W}}_{D}$. Using ${{\bm{W}}}_{U}{\bm{S}}_{k}$ and $\hat{{\bm{W}}}^{*}_{D}$ as the compressed matrices, the Type-I reconstruction error in equation 6 satisfies:
@@
where ${\bm{C}}_{\sigma}$ is the empirical correlation matrix of latent features ${\bm{C}}_{\sigma}=\sum_{i=1}^{N}\sigma({\bm{X}}_{i}{\bm{W}}_{U})^{\top}\sigma({\bm{X}}_{i}{\bm{W}}_{U})$, and (a) follows from Assumption 1. Setting the gradient of the last expression with respect to $\hat{{\bm{W}}}_{D}$ to zero, we obtain the optimal down matrix $\hat{{\bm{W}}}^{*}_{D}=\left({\bm{S}}_{k}^{\top}{\bm{C}}_{\sigma}{\bm{S}}_{k}\right)^{\dagger}{\bm{S}}_{k}^{\top}{\bm{C}}_{\sigma}{\bm{W}}_{D}$. After Plugging this back to the objective, we can further simplify the objective into,
$\min_{{\bm{S}}_{k}}\left\|\left({\bm{C}}^{\frac{1}{2}}_{\sigma}-{\bm{C}}^{\frac{1}{2}}_{\sigma}{\bm{S}}_{k}\left({\bm{S}}_{k}^{\top}{\bm{C}}_{\sigma}{\bm{S}}_{k}\right)^{\dagger}{\bm{S}}_{k}^{\top}{\bm{C}}_{\sigma}\right){\bm{W}}_{D}\right\|_{F}^{2}$
@@
Now, observe that the error on the right side of equation 14 is proportional to the Nyström matrix approximation to the matrix ${\bm{C}}_{\sigma}$ in Definition 1. Hence, the variable ${\bm{S}}_{k}$ can be optimized with any Nyström approximation algorithm [21]. In this work, we adapt a deterministic Nyström algorithm, Algorithm 1, that has the theoretical guarantee proved in the next proposition. ∎
@@
Suppose that the rank $k$ and the scores $s_{i}$ in Algorithm 1 are chosen such that there exists an error $\varepsilon>0$ satisfying $\varepsilon\geq\sum_{i=k+1}^{d_{int}}s_{i}$, then the Type-I modular reconstruction error in equation 6 is bounded by $V_{I}\leq\|{\bm{W}}_{D}\|_{2}^{2}\|{\bm{C}}^{-1}_{\sigma}\|_{2}\frac{\varepsilon^{2}{d^{2}_{int}}}{k^{2}(1-\varepsilon)^{2}}\sum_{i=k+1}^{d_{int}}\sigma_{i}^{2}({\bm{C}}_{\sigma})$, where $d_{\text{int}}$ and $\sigma_{i}$ denote the intermediate dimension (i.e., the input dimension of $W_{D}$) and singular values, respectively.
@@
Since our column selection is equivalent to applying the deterministic ridge leverage score sampling (DRLS) to ${\bm{C}}_{\sigma}^{\frac{1}{2}}$ [38] , Theorem 1 in [38] implies that
$(1-\varepsilon){\bm{C}}_{\sigma}-\frac{\epsilon}{k}\left\|({\bm{C}}_{\sigma}^{\frac{1}{2}})_{\backslash k}\right\|_{F}^{2}\mathbf{I}\preceq{\bm{C}}_{\sigma}^{\frac{1}{2}}{\bm{S}}_{k}{\bm{S}}_{k}^{\top}{\bm{C}}_{\sigma}^{\frac{1}{2}}\preceq{\bm{C}}_{\sigma}$ (15)
@@
Next, we define ${\bm{P}}={\bm{C}}^{\frac{1}{2}}_{\sigma}{\bm{S}}_{k}({\bm{S}}_{k}^{\top}{\bm{C}}_{\sigma}{\bm{S}}_{k})^{\dagger}{\bm{S}}_{k}^{\top}{\bm{C}}^{\frac{1}{2}}_{\sigma}$. We note that ${\bm{P}}$ is the projection matrix of the column space of ${\bm{C}}^{\frac{1}{2}}_{\sigma}{\bm{S}}$. Now, we multiply ${\bm{I}}-{\bm{P}}$ to both sides in the previous inequality to get
$({\bm{I}}-{\bm{P}}){\bm{C}}_{\sigma}({\bm{I}}-{\bm{P}})$ (17)
@@
where in the last inequality we use the fact that ${\bm{I}}-{\bm{P}}\preceq{\bm{I}}$ and that ${\bm{I}}-{\bm{P}}$ is the orthogonal projection to the orthogonal complement of the column space ${\bm{C}}_{\sigma}^{\frac{1}{2}}{\bm{S}}$ so that ${\bm{S}}^{\top}{\bm{C}}_{\sigma}^{\frac{1}{2}}({\bm{I}}-{\bm{P}})=\mathbf{0}$. Now, we have
$\|({\bm{I}}-{\bm{P}}){\bm{C}}^{\frac{1}{2}}_{\sigma}{\bm{C}}^{\frac{1}{2}}_{\sigma}({\bm{I}}-{\bm{P}})\|_{2}\leq\frac{\varepsilon}{k(1-\varepsilon)}\left\|({\bm{C}}_{\sigma}^{\frac{1}{2}})_{\backslash k}\right\|_{F}^{2}=\frac{\varepsilon}{k(1-\varepsilon)}\sum_{i=k+1}^{d_{\text{int}}}\sigma_{i}({\bm{C}}_{\sigma})$ (20)
@@
Since ${\bm{C}}^{\frac{1}{2}}_{\sigma}{\bm{P}}{\bm{C}}^{\frac{1}{2}}_{\sigma}={\bm{C}}_{\sigma}{\bm{S}}_{k}\left({\bm{S}}_{k}^{\top}{\bm{C}}_{\sigma}{\bm{S}}_{k}\right)^{\dagger}{\bm{S}}_{k}^{\top}{\bm{C}}_{\sigma}$, the inequality is equivalent to
$\|{\bm{C}}_{\sigma}-{\bm{C}}_{\sigma}{\bm{S}}\left({\bm{S}}^{\top}{\bm{C}}_{\sigma}{\bm{S}}\right)^{\dagger}{\bm{S}}^{\top}{\bm{C}}_{\sigma}\|_{2}\leq\frac{\varepsilon}{k(1-\varepsilon)}\sum_{i=k+1}^{d_{\text{int}}}\sigma_{i}({\bm{C}}_{\sigma}).$ (22)
@@
$\mathcal{E}_{\text{Nys}}^{2}({\bm{C}}_{\sigma})\leavevmode\nobreak$ $\overset{(a)}{\leq}{d_{\text{int}}}\|{\bm{C}}_{\sigma}-{\bm{C}}_{\sigma}{\bm{S}}\left({\bm{S}}^{\top}{\bm{C}}_{\sigma}{\bm{S}}\right)^{\dagger}{\bm{S}}^{\top}{\bm{C}}_{\sigma}\|^{2}_{2}$
$\overset{(b)}{\leq}\frac{\varepsilon^{2}{d_{int}}}{k^{2}(1-\varepsilon)^{2}}\left(\sum_{i=k+1}^{d_{int}}\sigma_{i}({\bm{C}}_{\sigma})\right)^{2}$
$\overset{(c)}{\leq}\frac{\varepsilon^{2}{d^{2}_{int}}}{k^{2}(1-\varepsilon)^{2}}\sum_{i=k+1}^{d_{int}}\sigma_{i}^{2}({\bm{C}}_{\sigma}),$ (23)